# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ 2 From Classical RL to CPT-RL
We will also be concerned with the particular case of (CPT-PO) in which $w^{+},w^{-}$ are set to the identity, namely the expected utility objective where only returns are distorted by the utility function:
@@
Similar problem variants for total cost and infinite horizon discounted settings can also be formulated. Notice that standard RL policy optimization problems are clearly particular cases of (CPT-PO).
Example: Personalized Treatment for Pain Management. We illustrate our problem formulation with a concrete example in healthcare to give the reader more intuition about the different features of CPT-RL, its importance in applications when human perception and behavior matter and its differences compared to risk-sensitive RL. The goal is to help a physician manage a patient’s chronic pain by suggesting a personalized treatment plan over time. The challenge here is to balance pain relief and the risk of opioid dependency or other side effects that might be due to the treatment, i.e. short-term relief and longer term risks. We propose to train a CPT-RL agent to help the physician.
  1. 1.
    Why sequential decision making? (a) The physician needs to adjust treatment at each time step depending on the patient’s reported pain level as well as the observed side effects. This is relevant to dynamic treatment regimes in general (such as for chronic diseases, see e.g. Yu et al. (2021) for a survey) in which considering delayed effects of treatments is also important (and RL does account for such effects). (b) Decisions clearly impact the patient’s immediate pain relief, dependency risks in the future and their overall health condition. A state is described by e.g. current pain level, dependency risk and side effect severity. Actions are treatments, e.g. no treatment, alternative treatment or opioid treatment.
  2. 2.
    Why CPT-RL? Patients and clinicians make decisions influenced by psychological biases. We illustrate the importance of each one of the three features of CPT in section 2 (reference point, utility and probability distortion weight functions) via this example:
    • Utility transformation: Patients might often show a loss averse behavior, i.e., they might perceive pain increase or withdrawal symptoms as worse than equivalent gains in pain relief. Note here that loss aversion should not be confused with risk aversion (Schmidt and Zank, 2005). In short, loss aversion can be defined as a cognitive bias in which the emotional impact of a loss is more intense than the satisfaction derived from an equivalent gain. For instance, in our example, a 2-point increase in pain might be seen as much worse than a 2-point reduction even if the change is the same in absolute value. This loss aversion concept is a cornerstone of Kahneman and Tversky’s theory. In contrast, risk aversion rather refers to the rational behavior of undervaluing an uncertain outcome compared to its expected value. Risk sensitive approaches might be less adaptive to a patient’s subjective preferences if they deviate from objective risk assessments.
  3. 3.
    CPT-RL vs Risk-averse RL. In terms of policies, risk-averse RL would favor non-opioid treatments unless extreme pain levels make opioids justifiable. In contrast, CPT-RL policies would prescribe opioids if pain significantly exceeds the patient’s reference point. As dependency risk increases, CPT-RL policies would transition to non-opioid treatments as a consequence of overweighting the probability of rare catastrophic outcomes. Notably, CPT-RL policies can oscillate between risk-seeking (to address high pain) and risk-averse (to avoid severe side effects). In contrast, a risk-sensitive agent focuses on minimizing variability in health states and dependency risks and would likely avoid opioids in most cases unless pain levels become extreme. Such risk-sensitive policies favor stable strategies (e.g., consistent non-opioid use), prioritizing low variance in patient outcomes.
Scope of this work. Several behavioral studies model the reference points, utility and probability weight functions using real patient treatment data (Mkrtchian et al., 2023; George et al., 2019; Sip et al., 2018). In real-world applications, known utility and probability weighting functions could originate from domain experts, surveys, or prior empirical studies on target user groups (see appendix F.6 for more details on the choice of the utility function) . In this work, we suppose that these are known. Our goal is to align the agent’s behavior with the given preferences by optimizing for the CPT value of returns. The known functions can be interpreted as the mathematical representation of the preferences that the agent must align with. Our main challenge consists in ensuring policy alignment with those predefined human preferences. We leave the question of discovering or inferring human preferences (e.g., inverse RL, RL from Human Feedback (RLHF)) which is central in many alignment problems to future work. See section E for an extended discussion comparing CPT-RL to RLHF as preference learning paradigms, their pros and cons and opportunities for future work in combining them as they are not mutually exclusive. We also refer the reader to appendix D for an extended discussion regarding applications.
@@
Challenges. To conclude this section, we describe the challenges we face in solving (CPT-PO). First, the CPT value does not satisfy a Bellman equation due to the nonlinearity of the utility and weight functions which breaks the additivity and linearity of the standard expected return. Second, (CPT-PO) is a nonconvex problem involving several nonconvex functions: The utility itself is nonconvex in general (recall the utility is convex w.r.t. gains and concave w.r.t. losses) and the probabilities are also distorted by a nonconvex weight function. While the standard policy optimization problem is already nonconvex in the policy, (CPT-PO) further introduces additional nonconvexity.
3 About Optimal Policies in CPT-RL
In this section, we investigate the properties of optimal policies to (CPT-PO) when they exist. We focus on constrasting our results with existing known results for solving standard MDPs to highlight the peculiarities of our CPT-RL problem. Understanding the properties of optimal policies are important in view of designing efficient policy search algorithms.
We start our discussion by pointing out a stark difference between optimal policies in standard MDPs and (CPT-PO). While there exists an optimal deterministic stationary policy for MDPs (see e.g. Thm. 6.2.10 in Puterman (2014)), this is not the case in general for (CPT-PO).
@@ 3 About Optimal Policies in CPT-RL
There does not always exist an optimal policy for (CPT-PO) in $\Pi_{NM}^{D}$ (i.e. deterministic non-Markovian).
Proposition 2 tells us that the stochasticity of the policy is essential in solving our CPT-RL problem. The proof of this result is deferred to Appendix G.2: we construct a simple problem instance where an optimal policy needs to be stochastic as any deterministic policy is necessarily and clearly suboptimal. Our example is built around a $w^{+}$ function that puts special emphasis on the $10\%$ of the best outcomes. As a consequence, the optimal policy needs to be randomized to take advantage of this and obtain the highest returns with some probability without suffering from bad outcomes by deterministically committing to this riskier strategy. It has been briefly mentioned in L.A. et al. (2016) that the policy needs to be random in general for (CPT-PO), see also the organ transplant example in Lin et al. (2018).
The next result shows that the need for stochasticity in the optimal policy is clearly due to the probability distortions in the CPT value. Indeed, when setting the probability weight distortion function $w$ to the identity, i.e. when considering the particular case (EUT-PO) of (CPT-PO), it appears that an optimal policy is not necessarily stochastic.
Proposition 3.
There exists an optimal policy for (EUT-PO) in $\Pi_{\Sigma,NS}^{D}\,.$
Proposition 3 allows to safely restrict our policy search to $\Pi_{\Sigma,NS}$ which is a much smaller policy space than the set of non-Markovian policies $\Pi_{NM}$. The fact that an optimal deterministic policy exists is a fundamental difference with the general (CPT-PO) setting. Whether there are specific weight functions (apart from the identity) for which there always exist a deterministic optimal policy remains an open question that we leave for future work.
Proposition 3 also shows that (EUT-PO) is simpler than the more general single-trial RL problem (Mutti et al., 2023a) in which one needs to look for an optimal policy in a much larger policy set $\Pi_{NM}^{D}$ than $\Pi_{\Sigma,NS}^{D}$ in general. See appendix F.5 for the connection between both problems.
We now ask the next natural question: Can we further restrict our policy search to a smaller policy class compared to $\Pi_{\Sigma,NS}\,$? In particular, are there specific utility functions for which the resulting (EUT-PO) problem has optimal Markovian policies? We provide a positive answer by establishing a precise characterization of such utility functions which turn out to be either affine or exponential.
Theorem 4.
Let $\mathcal{U}$ be continuous and increasing. The following statements are equivalent: 1. For any MDP, there exists an optimal policy for (EUT-PO) in $\Pi_{M,NS}$ 2. There exists $\varphi:\mathbb{R}^{2}\to\mathbb{R}$ s.t.$\forall x,a,b\in\mathbb{R},b\neq 0,$
$\mathcal{U}(x+a)-\mathcal{U}(x)=\varphi(a,b)(\mathcal{U}(x+b)-\mathcal{U}(x))\,.$
3. There exists $\mu:\mathbb{R}^{2}\to\mathbb{R}$ s.t.$\forall y,c,d\in\mathbb{R},$
$\mathcal{U}(y+c)-\mathcal{U}(c)=\mu(c,d)(\mathcal{U}(y+d)-\mathcal{U}(d))\,.$
4. There exist $A,B,C\in\mathbb{R}$ s.t. $\mathcal{U}(x)=Ax+B$ or $\mathcal{U}(x)=A+B\exp(Cx)$ for all $x\in\mathbb{R}\,.$
A few comments are in order regarding Theorem 4:
(i) So far, we have highlighted the importance of the probability distortion function in determining the nature of optimal policies for (CPT-PO). Theorem 4 is rather concerned with the role of the (nonlinear) utility functions in (CPT-PO).
(ii) The theorem is reminiscent of the following known folklore result: The only memoryless continuous probability distribution is the exponential distribution.
(iii) Theorem 4 shows that the only utility functions leading to optimal Markovian policies are the affine and exponential utilities. The affine utility makes (CPT-PO) boil down to a standard RL problem whereas the exponential criterion is a well-known objective used in the risk-sensitive control and RL (see section 6 and appendix C for a discussion).
(iv) Monotonicity of the utility function is a fundamental assumption in economics and decision theory (utility increases with increasing gains). It is not crucial for most of our later developments though, see appendix F.7 for a discussion.
(v) Recall that our result holds for the finite horizon undiscounted setting which is not restrictive in applications. The discounted infinite horizon setting introduces additional technical challenges due to the lack of positive homogeneity of the CPT value, see Hau et al. (2023) (and Theorem 3.3 therein) for a discussion of this question for the particular case of the entropic risk measure. The use of positive quasi-homogeneity might allow to extend our result using similar techniques. We leave this for future work.
Theorem 4 is concerned with (EUT-PO) which is a particular case of (CPT-PO). However, these results cannot be extended to (CPT-PO) in general as we show next.
Proposition 5.
@@
There exist instances of (CPT-PO) where $\mathcal{U}$ is of the form $x\mapsto A+B\exp(Cx)$ for positive constants $A,B,C$ and (CPT-PO) does not admit an optimal policy in $\Pi_{M,NS}\,.$
4 Policy Gradient Algorithm for CPT-value Maximization
@@
In this section, we propose a policy gradient algorithm for solving (CPT-PO). From this section on, we parametrize policies $\pi\in\Pi_{NM}$ by a vector $\theta\in\mathbb{R}^{d}$ and we denote by $\pi_{\theta}$ the parametrized policy. As a consequence, the CPT objective in (CPT-PO) becomes a function of the policy parameter $\theta$ and we use the shorthand notation $J(\theta)$ for the corresponding CPT objective value.
@@
[…]
@@
Stochastic Policy Gradient Algorithm for CPT-RL. In the light of Theorem 6, we will perform a policy gradient ascent on the objective $J$ to solve (CPT-PO). Our general policy gradient algorithm is presented in Algorithm 1. As usual, since we only have access to sampled trajectories from the MDP, we need a stochastic policy gradient to estimate the true unknown gradient given by the theorem. In particular, we need an approximation of $\varphi(R(\tau))$ for any sampled trajectory $\tau$ from the MDP following policy $\pi_{\theta}$. In the case of (EUT-PO) in which $w$ is the identity, the unknown quantity $\varphi(R(\tau))$ reduces to $\mathcal{U}(R(\tau))$ which can be easily computed as $\mathcal{U}$ is known and $R(\tau)$ is the cumulative reward.
Algorithm 1 CPT-Policy Gradient Algorithm (CPT-PG)
1:Input: $\theta_{0}\in\mathbb{R}^{d}$, utility functions $u^{+},u^{-}$, weight functions $w_{+},w_{-}$, step size $\alpha>0\,.$
@@
4:     Sample a trajectory $\tau:=(s_{t},a_{t},r_{t})_{0\leq t\leq H-1}$, with $s_{0}\sim\rho$ following $\pi_{\theta_{k}}$
@@
10:     $\hat{\phi}_{n}^{\pm}=\sum_{i=0}^{j_{n}-1}w^{\prime}_{\pm}\left(\frac{i}{n}\right)\left(\hat{\xi}^{\pm}_{\frac{n-i}{n}}-\hat{\xi}^{\pm}_{\frac{n-i-1}{n}}\right)+w^{\prime}_{\pm}\left(\frac{j_{n}}{n}\right)\left(R(\tau)-\hat{\xi}^{\pm}_{\frac{n-j_{n}-1}{n}}\right)$
11:     $\hat{g}_{k}=(\hat{\phi}_{n}^{+}-\hat{\phi}_{n}^{-})\sum_{t=0}^{H-1}\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}|h_{t})\,$
12:/Policy gradient update
13:     $\theta_{k+1}=\theta_{k}+\alpha\,\hat{g}_{k}$
14:end for
In the more general setting, the approximation task requires to compute the term $\int w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z)dz$ (and likewise for the second integral term). We address this challenge using the following result which is a variation of Proposition 6 in L.A. et al. (2016) in which the integrand is the derivative $w_{+}^{\prime}$ (instead of $w_{+}$) and the integral is taken over a bounded interval. While L.A. et al. (2016) use this result to approximate the CPT value, we use it for approximating our special integral terms involving the derivatives of the weight functions as they appear in the policy gradient. We obtain a different approximation formula which is tailored to our setting. The approximation is essentially a Riemann sum using simple staircase functions.
@@
Similarly to L.A. et al. (2016) (Theorem 1), our algorithm can be shown to enjoy a similar asymptotic convergence result to the set of stationary points of the (CPT-PO) objective. This is because we can also employ the same stochastic approximation artillery upon noticing that we are also approximating the same policy gradient differently and the induced bias in our case will also vanish with a large enough number of trajectories $n$ (by Thm. 6 and Prop. 16). Notice that we can also remove the projection therein upon assuming that the rewards and the score function in the policy gradient are both bounded. These fairly standard assumptions in the analysis of vanilla PG methods guarantee that the policy gradient will remain bounded.
@@
Comparison to the CPT-SPSA-G algorithm in L.A. et al. (2016). Our algorithm is specifically designed for maximizing the CPT value of a sum of rewards generated by an MDP while the CPT-SPSA-G algorithm in L.A. et al. (2016) can be used for a larger class of problems to maximize the CPT value of any real-valued random variable. However, we highlight that (a) this cumulative reward return structure is natural and ubiquitous in RL and economics applications and foremost (b) thanks to this particular problem structure, our algorithm is a policy gradient algorithm leveraging first-order information whereas CPT-SPSA-G only uses zeroth order information, i.e. CPT value estimations. This difference is crucial as zeroth order optimization algorithms are known to suffer from the curse of dimensionality. Our algorithm can scale better to higher dimensional problems as it is notoriously known for policy gradient algorithms in classic RL. We provide empirical evidence of this fact in section 5 to further support the benefits of our algorithm.

### v3

@@ 2.2 CPT-RL Problem Formulation
For an illustration of the CPT-RL problem formulation and its elements in a personal treatment for pain management application, see App. E.5. In the particular case of CPT-PO in which $w^{+},w^{-}$ are set to the identity, namely the Expected Utility Policy Optimization (EUT-PO) objective, only returns are distorted by the utility function:
@@
Challenges. We outline the main difficulties in solving CPT-PO. First, the CPT value does not satisfy a Bellman equation: the nonlinear utility and weight functions violate the additivity and linearity of the standard expected return. While the special case EUT-PO has been studied (see e.g. Bäuerle and Rieder (2014); Fei et al. (2020); Wu and Xu (2023)), the CPT setting introduces fundamental differences. In particular, probability distortion breaks the dynamic programming structure, and optimal policies may be stochastic even under identity utility. Crucially, this aspect, central to CPT-RL and enabling richer behavioral modeling, has not been addressed in prior work on EUT-PO. Second, CPT-PO is highly nonconvex. The utility function is nonconvex in general (convex over gains, concave over losses), and the probabilities are also distorted by a nonconvex weight function. While the standard policy optimization problem is already nonconvex, CPT-PO compounds this difficulty with additional sources of nonconvexity.
Scope of this work. We assume the utility and weighting functions are known, e.g., from domain experts, surveys, prior empirical studies on target user groups, or behavioral studies (Mkrtchian et al., 2023; George et al., 2019; Sip et al., 2018), see App. G.7 for more details on the choice of the utility function. Our goal is to align the agent’s behavior with the given preferences (encoded in utility and weighting functions) by optimizing for the CPT value of returns.
We leave the question of discovering or inferring human preferences (e.g. RL from Human Feedback (RLHF)) to future work. See section F for an extended discussion comparing CPT-RL to RLHF as preference learning paradigms, their pros and cons and opportunities for future work in combining them as they are not mutually exclusive. See also App. E for applications.
2.3 Peculiarities of Optimal Policies in CPT-RL
In this section, we highlight the properties of optimal policies to CPT-PO when they exist and contrast them with classical RL and EUT-PO (details in App. H):
The need for stochastic policies. In MDPs, optimal deterministic stationary policies always exist; in CPT-RL, optimal policies may need to be stochastic. In general, stochasticity of the policy is essential in solving CPT-PO. See App. H.1 for a proof.
Importance of probability weighting. Under EUT-PO, deterministic non-Markovian policies suffice (see e.g. Theorem 1 in Bäuerle and Rieder (2014)), but this is not the case for CPT-PO with probability weighting in general. Therefore, the need for stochasticity in the optimal policy is clearly due to the probability distortions in the CPT value.
The need for non-Markovian policies. Under EUT-PO, optimal Markovian policies exist for special cases of utility functions. Bäuerle and Rieder (2014) establish a characterization of such utility functions which turn out to be either affine or exponential (when $\mathcal{U}$ is continuous and increasing). This highlights the role of the (nonlinear) utility functions on the nature of optimal policies. However, this result cannot be extended to CPT-PO in general as we show next.
@@ 2.3 Peculiarities of Optimal Policies in CPT-RL
There exist instances of CPT-PO where $\mathcal{U}$ is of the form $x\mapsto A+B\exp(Cx)$ for positive constants $A,B,C$ and CPT-PO does not admit an optimal policy in $\Pi_{M,NS}\,.$
Even exponential utilities, which guarantee the existence of optimal Markovian policies in EUT-PO, may fail to admit Markovian optimal policies in CPT-PO.
3 Policy Gradient for CPT-Value Maximization
@@
In this section, we design a policy gradient algorithm for solving CPT-PO. From this section on, we parametrize policies $\pi\in\Pi_{NM}$ by a vector $\theta\in\mathbb{R}^{d}$ and we denote by $\pi_{\theta}$ the parametrized policy. As a consequence, the CPT objective in CPT-PO becomes a function of the policy parameter $\theta$ and we use the notation $J(\theta)$ for the corresponding CPT objective value.
3.1 CPT Policy Gradient Theorem
@@
(Policy Gradient for CPT-RL). Suppose that the utility functions $u^{-},u^{+}$ are continuous and that the weight functions $w_{-},w_{+}$ are Lipschitz and differentiable. Assume in addition that the policy parametrization $\theta\mapsto\pi_{\theta}(a|h)$ (for any $h,a\in\mathcal{H}\times\mathcal{A}$) are both differentiable. Then, for every $\theta\in\mathbb{R}^{d}$, the gradient of the CPT-PO objective $J$ w.r.t. the policy parameter $\theta$ is given by:
$\nabla J(\theta)=\mathbb{E}\left[\varphi\left(R(\tau)\right)\sum_{t=0}^{H-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}|h_{t})\right]\,$,
where $R(\tau):=\sum_{t=0}^{H-1}r_{t}\,$ with $\tau:=(s_{t},a_{t},r_{t})_{0\leq t\leq H-1}$ is a trajectory of length $H$ generated from the MDP by following policy $\pi_{\theta}\,$ [^3] The integral $\varphi(R(\tau))$ is finite under our continuity assumptions since the return $R(\tau)$ is bounded. and for all $v\in\mathbb{R},\varphi(v):=\int_{z=0}^{\max(v,0)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz$ $\int_{z=0}^{\max(-v,0)}w_{-}^{\prime}(\mathbb{P}(u^{-}((R(\tau)>z))dz,$ and $w_{+}^{\prime},w_{-}^{\prime}$ denoting the derivatives.
@@
3.2 Stochastic PG Algorithm for CPT-RL
In the light of Theorem 3, we will perform a policy gradient ascent on the objective $J$ to solve CPT-PO. Our general PG algorithm is presented in Algorithm 1. As usual, since we only have access to sampled trajectories from the MDP, we need a stochastic policy gradient to estimate the true unknown gradient given by the theorem. In particular, we need an approximation of $\varphi(R(\tau))$ for any sampled trajectory $\tau$ from the MDP following policy $\pi_{\theta}$. In the case of EUT-PO in which $w$ is the identity, the unknown quantity $\varphi(R(\tau))$ reduces to $\mathcal{U}(R(\tau))$ which can be easily computed as $\mathcal{U}$ is known and $R(\tau)$ is the cumulative reward.
Algorithm 1 CPT-Policy Gradient (CPT-PG)
1:Input: $\theta_{0}\in\mathbb{R}^{d}$, utility functions $u^{+},u^{-}$, weight functions $w_{+},w_{-}$, step sizes $(\alpha_{k})\,.$
@@ 3.2 Stochastic PG Algorithm for CPT-RL
4:  Sample $m$ trajectories $\tau_{l}:=(s_{t}^{l},a_{t}^{l},r_{t}^{l})_{0\leq t\leq H-1}$, $1\leq l\leq m$ with $s_{0}^{l}\sim\rho$ following $\pi_{\theta_{k}}$
@@
10:  $\hat{\phi}_{k,n}^{\pm}=\sum_{i=0}^{j_{n}-1}w^{\prime}_{\pm}\left(\frac{i}{n}\right)\left(\hat{\xi}^{\pm}_{\frac{n-i}{n}}-\hat{\xi}^{\pm}_{\frac{n-i-1}{n}}\right)+w^{\prime}_{\pm}\left(\frac{j_{n}}{n}\right)\left(R(\tau)-\hat{\xi}^{\pm}_{\frac{n-j_{n}-1}{n}}\right)$
11:  $\psi_{k,n}=\frac{1}{m}\sum_{l=1}^{m}\sum_{t=0}^{H-1}\nabla_{\theta}\log\pi_{\theta_{k}}(a_{t}^{l}|h_{t}^{l})$
12:  $\hat{\nabla}_{n,m}J(\theta_{k})=(\hat{\phi}_{n}^{+}-\hat{\phi}_{n}^{-})\cdot\psi_{k,n}\,$
13:/Policy gradient update
14:  $\theta_{k+1}=\theta_{k}+\alpha_{k}\,\hat{\nabla}_{n,m}J(\theta_{k})$
15:end for
In the more general setting, the approximation task requires to compute the integral term $\int w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z)dz$ (and likewise for the second integral term). We address this challenge using the following result which is a variation of Proposition 6 in L.A. et al. (2016) in which the integrand is the derivative $w_{+}^{\prime}$ (instead of $w_{+}$) and the integral is taken over a bounded interval. While L.A. et al. (2016) use this result to approximate the CPT value, we use it for approximating our special integral terms involving the derivatives of the weight functions as they appear in the policy gradient. We obtain a different approximation formula which is tailored to our setting. The approximation is essentially a Riemann sum using simple staircase functions.
@@
Let $X$ be a real-valued random variable. Suppose that the functions $w^{\prime}_{+},w^{\prime}_{-}$ are Lipschitz and that $u^{+}(X),u^{-}(X)$ have bounded first moments. Let $\xi^{+}_{\frac{i}{n}}$ and $\xi^{-}_{\frac{i}{n}}$ denote the $\frac{i}{n}$th quantile of $u^{+}(X)$ and $u^{-}(X)$, respectively. Then, we have for any $v\geq 0$,
$\int_{0}^{v}w^{\prime}_{+}(\mathbb{P}(u^{+}(X)>z))dz=\lim_{n\rightarrow\infty}u_{n}\,$,
@@
where $j_{n}\in[0,n-1]$ is s.t. $v\in[\xi^{+}_{\frac{n-j_{n}-1}{n}},\xi^{+}_{\frac{n-j_{n}}{n}}$]\,. The same identity holds when replacing $u^{+}(X),\xi^{+}_{\alpha},w_{+}$ by $u^{-}(X),\xi^{-}_{\alpha},w_{-}$ where $\xi^{-}_{\alpha}$ is the $\alpha^{\text{th}}$ quantile of $u^{-}(X)\,.$
@@
Comparison to the CPT-SPSA-G algorithm. Our algorithm is designed for maximizing the CPT value of a sum of rewards generated by an MDP while the CPT Simultaneous Perturbation Stochastic Approximation Gradient (CPT-SPSA-G) algorithm in L.A. et al. (2016) can be used to maximize the CPT value of any real-valued random variable. However, we highlight that (a) this cumulative reward return structure is natural and ubiquitous in RL and economics applications and foremost (b) thanks to this problem structure, our PG algorithm leverages first-order information whereas CPT-SPSA-G only uses zeroth-order information, i.e. CPT value estimations. This difference is crucial as zeroth-order optimization algorithms are known to suffer from the curse of dimensionality. Our algorithm can scale better to higher dimensional problems as it is notoriously known for PG algorithms in classic RL. We provide empirical evidence of this fact in section 5 to further support the benefits of our algorithm.
4 Asymptotic Convergence and Sample Complexity
In this section, we establish asymptotic convergence and sample complexity guarantees for both our PG estimator and Algorithm 1.
First, we show that our PG estimator is consistent.
Proposition 5 (Consistency).
Suppose that the utility functions $u^{+}$ and $u^{-}$ are continuous, strictly increasing and uniformly bounded by a positive constant $M_{u}\,.$ Assume in addition that the functions $w^{\prime}_{+},w^{\prime}_{-}$ are $L_{w}$-Lipschitz. Then for any $\theta\in\mathbb{R}^{d},\hat{\nabla}_{n,m}J(\theta)\to\nabla J(\theta)$ almost surely as the batch size parameters $n,m\to\infty\,.$
To prove this result, we combine consistency of the Monte Carlo estimators of the expected score function ($\nabla\ln\pi_{\theta}$ sum term) and the expected distorted reward ($\phi(R(\tau))$ term) in Thm. 3, using independent random trajectories. Consistency of the first follows from the law of large numbers whereas the second follows from using a generalized dominated convergence theorem combined with the Glivenko-Cantelli theorem (to control the asymptotics of the empirical distribution function stemming from quantile estimation), similarly to the proof of Prop. 3 in L.A. et al. (2016).
Beyond consistency, we now quantify the number of trajectories ($n+m$) required to obtain an $\varepsilon$-approximate policy gradient for any policy parameter. Under the same assumptions as for the consistency result, we make an additional score boundedness assumption which is standard in the analysis of PG methods (see e.g. Papini et al. (2018); Yuan et al. (2022); Fatkhullin et al. (2023)).
Proposition 6 (Sample complexity).
Let the assumptions of Prop. 5 hold. Suppose in addition that the score function is bounded, i.e. there exists $M_{\psi}>0$ s.t. $\|\nabla\ln\pi_{\theta}(a|s)\|_{2}\leq M_{\psi}$ for all $\theta\in\mathbb{R}^{d},(s,a)\in\mathcal{S}\times\mathcal{A}$. Then there exists an absolute constant $c>0$ s.t. for any $\varepsilon>0,\delta\in(0,1)$ and for all $\theta\in\mathbb{R}^{d},$ we have $\|\hat{\nabla}_{n,m}J(\theta)-\nabla J(\theta)\|_{2}\leq\varepsilon$ with probability at least $1-\delta$, if $n\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(1/\delta)}{\varepsilon^{2}}$ and $m\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(2d/\delta)}{\varepsilon^{2}}\,.$
The sample complexity result coincides with the classical $n+m=\tilde{\mathcal{O}}(\varepsilon^{-2})$ statistical rate of Monte Carlo estimation. The sample complexity increases with the curvature of the probability weight function, the magnitude of the utility function and the horizon length. The proof relies on using concentration inequality results, namely (i) a Hoeffding’s style inequality for bounded vector-valued random variables (Jin et al., 2019) for the score function and (ii) Dvoretzky-Kiefer-Wolfowitz inequality to quantify concentration of the empirical distribution of a random variable. This second part of the proof follows similar lines as the proof of Prop. 3 in L.A. et al. (2016) where it rather used for CPT value estimation rather than our distorted reward (see Thm. 3 and Prop. 4). Note though that the direct dependence on the dimension $d$ of the policy parameter is only logarithmic in contrast to the sample complexity of zeroth-order PG estimation which scales with the dimension $d$ in L.A. et al. (2016) due to the need to estimate each policy gradient coordinate.
We close this section by discussing the convergence of the policy parameter sequence produced by Algorithm 1. The CPT-RL objective is non-convex in the policy parameter due to non-convexity of both utility and probability weighting functions. While this lack of structure makes global optimality out of reach, we can still obtain a standard asymptotic convergence result towards the set of stationary points using the machinery of stochastic approximation (see e.g. (Borkar and Borkar, 2008; Benaïm, 2006)).
Proposition 7 (Asymptotic convergence).
Under the same assumptions as in Prop. 5, suppose in addition that the positive step sizes $(\alpha_{k})$ satisfy the Robbins-Monro conditions $\sum_{k}\alpha_{k}=+\infty$ and $\sum_{k}\alpha_{k}^{2}<+\infty\,.$ Then the sequence of iterates $(\theta_{k})$ generated by Algorithm 1 converges to the set of stationary points of $J$ almost surely, i.e. $\theta_{k}\to\{\theta\in\mathbb{R}^{d}:\nabla J(\theta)=0\}$ almost surely as $k\to\infty$ if the sequence $(\theta_{k})$ is bounded.
Compared to the asymptotic convergence result established for the zeroth-order CPT-SPSA-G algorithm proposed in L.A. et al. (2016), note that we do not use increasing batch sizes to reduce the bias due to zeroth order PG estimation as our PG estimator is unbiased. Our boundedness assumption can be simply relaxed by adding a projection in the gradient ascent step in Algorithm 1 and modifying the limit set to account for the projection similarly to the statement of Thm. 1 in L.A. et al. (2016). We prefer our simpler statement.